# Randomized Black-Box PIT for Small Depth +-Regular Non-commutative Circuits

### v1

@@ Abstract
In particular, we show that if $f$ is a non-zero non-commutative polynomial in $n$ variables over the field $\mathbb{F}$, computed by a depth-$d$ $+$-regular circuit of size $s$, then $f$ cannot be a polynomial identity for the matrix algebra $\mathbb{M}_{N}(\mathbb{F})$, where $N=s^{O(d^{2})}$ and the size of the field $\mathbb{F}$ depending on the degree of $f$. Our result can be interpreted as an Amitsur-Levitzki-type result [AL50] for polynomials computed by small-depth $+$-regular circuits.
@@ 1 Introduction
Recently, [AJMR19] gave an efficient randomized algorithm for the PIT problem when the circuits are allowed to compute polynomials of exponential degree, but the sparsity could be exponential in the size of the circuit. To handle doubly-exponential sparsity, they studied a class of homogeneous non-commutative circuits, that they call +-regular circuits, and gave an efficient deterministic white-box PIT algorithm. These circuits can compute non-commutative polynomials with the number of monomials doubly exponential in the circuit size. For the black-box setting, they obtain an efficient randomized PIT algorithm only for depth-3 +-regular circuits. In particular, they show that if a non-zero non-commutative polynomial $f\in\mathbb{F}\mathopen{\langle}X\mathclose{\rangle}$ is computed by a depth-3 +-regular circuit of size $s$, then $f$ cannot be a polynomial identity for the matrix algebra $\mathbb{M}_{s}(\mathbb{F})$ for a sufficiently large field $\mathbb{F}$. Finding an efficient randomized PIT algorithm for higher depth +-regular circuits is listed as an interesting open problem. We resolve this problem for constant depth +-regular circuits. In particular, we show that if $f\in\mathbb{F}\mathopen{\langle}X\mathclose{\rangle}$ is a non-zero non-commutative polynomial computed by a depth-$d$ +-regular circuit of size $s$, then $f$ cannot be a polynomial identity for the matrix algebra $\mathbb{M}_{N}(\mathbb{F})$, with $N=s^{O(d^{2})}$ and the size of the field $\mathbb{F}$ depends on the degree of polynomial $f$. This resolves an open problem given in [AJMR19]. We note that we get a black-box randomized polynomial time black-box PIT algorithm for constant depth $+$-regular circuits.
@@
The size of the $+$-regular circuit is the number of gates in the circuit. As noted earlier, the non-commutative polynomial $(x+y)^{2^{s}}$ can be computed by a depth-3 $+$-regular circuit, denoted by $\Sigma\Pi^{*}\Sigma$, of size $O(s)$ using repeated squaring. This circuit consists of two addition layers, namely $\mathcal{L}_{1}^{+},\mathcal{L}_{2}^{+}$ and two multiplication layers, namely $\mathcal{L}_{1}^{\times},\mathcal{L}_{2}^{\times}$. The multiplication layer $\mathcal{L}_{1}^{\times}$ consists of only the two input gates labeled by $x$ and $y$ respectively. The addition layer $\mathcal{L}_{1}^{+}$ consists of only one addition gate computing the homogeneous linear form $(x+y)$. The multiplication layer $\mathcal{L}_{2}^{\times}$ consists of only one gate computing the polynomial $(x+y)^{2^{s}}$. The addition layer $\mathcal{L}_{2}^{+}$ consists of only the ouptut gate computing the polynomial $(x+y)^{2^{s}}$. In this example, the top-most addition gate (i.e., the output gate) essentially has one input. The main result of the paper is the following theorem.
@@
For degree $D$ non-zero non-commutative polynomial $f$, the classical Amitsur-Levitzki [AL50] theorem guarantees that $f$ does not vanish on the matrix algebra $\mathcal{M}_{\frac{D}{2}+1}(\mathbb{F})$. If $D=2^{\Omega(s)}$, this gives us an exponential time randomized PIT algorithm [BW05], where $s$ is the size of the circuit computing $f$. If the sparsity of the polynomial, i.e., the number of non-zero monomials, is doubly exponential, then the main result of [AJMR19] gives only an exponential time randomized PIT algorithm as their matrix dimension depends on the logarithm of the sparsity.
@@
We note that [AJMR19] presented a white-box deterministic polynomial-time PIT for arbitrary depth $+$-regular circuits. For the small-degree case, [RS05] provided a white-box deterministic polynomial-time PIT for non-commutative ABPs, while [FS13, AGKS15] have shown a quasi-polynomial-time black-box PIT algorithm for non-commutative ABPs.
@@ 1.1 Outline of the Proofs: High-level Idea
In the rest of the paper, we will use ”n.c.” as an abbreviation for ”non-commutative”. We first explain the main ideas behind the randomized PIT algorithm for depth-5 $+$-regular circuits, as this is the major bottleneck. Consider an n.c. polynomial $f$ over $X=\{x_{1},\ldots,x_{n}\}$ computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$. The polynomial $f$, computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$, can be written as follows:
@@ 1.1.1 Step 1: Transforming the Polynomial for Improved Structure
This fact is formalized in Lemma 4.1.1. However, this fact alone will not be sufficient to design an identity-testing algorithm for depth-5 +-regular circuits. This is because, to choose a small index set from each $Q_{ij}$, one would have to know the boundary between $Q_{ij}$ and $Q_{i,j+1}$ (i.e., the position at which $Q_{ij}$ ends and $Q_{i,j+1}$ begins), for any $i\in[s],j\in[D_{2}-1]$. To know where $Q_{ij}$ ends and $Q_{i,j+1}$ begins exactly, the algorithm will have to keep track of the degree (length) of $Q_{ij}$. This is not feasible as the degree of $Q_{ij}$ could be exponential in $s$. This is one of the main challenges in designing a black-box identity-testing algorithm for depth-5 +-regular circuits and is one of the main differences between depth-3 and depth-5 +-regular circuits (small depth $+$-regular circuits, in general).
@@ 2.1 Substitution Automaton
Now, we explain how a substitution automaton $\mathcal{A}$ processes a given polynomial. First, we will describe the output of a substitution automaton $\mathcal{A}$ on a monomial $w\in X^{D}$. Let $s$ and $t$ be the designated initial and final states of $\mathcal{A}$, respectively. For each variable $x\in X$, we have a transition matrix $M^{\prime}_{x}$ of size $|Q|\times|Q|$. If $w=x_{j_{1}}x_{j_{2}}\ldots x_{j_{d}}$, then the output of the substitution automaton $\mathcal{A}$ on the monomial $w$ is given by the $(s,t)$-th entry of the matrix $M^{\prime}_{w}=M^{\prime}_{x_{j_{1}}}M^{\prime}_{x_{j_{2}}}\ldots M^{\prime}_{x_{j_{D}}}$. Note that this entry can be a polynomial since there could be multiple paths from $s$ to $t$ labeled by the same monomial $w$.
@@ 4 Black-Box PIT for $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ Circuits
Definition 4.0.2 ($\xi$-Pattern).
We refer to the string $\xi_{1}^{\ell_{1}}\xi_{2}^{\ell_{2}}\cdots\xi_{s}^{\ell_{s}}$ as a $\xi$-pattern, regardless of the specific exponents of the $\xi$ variables.
@@ 4.1.1 Idea of the Substitution Automaton
    As explained in the proof outline, one would have to identify the boundary between $Q_{ij}$ and $Q_{i,j+1}$ (i.e., the position at which $Q_{ij}$ ends and $Q_{i,j+1}$ begins), for any $i\in[s],j\in[D_{2}-1]$.
@@
    We need to ensure that the resulting commutative polynomial is non-zero because the same commutative variables appearing in different transformed $Q_{ij}$ polynomials get mixed (i.e., exponents of the same variable appearing at different transformed $Q_{ij}$ polynomials are added).
@@
which is a polynomial in the variables $\xi\sqcup Y\sqcup Z$. This differs from previous works ([AJMR19, BW05]) where only a single matrix entry was used as the output.
@@
Next, we consider the output of the substitution automaton $\mathcal{A}$ on $m$. The automaton knows how to replace any variable $x_{j}$ at any state $q$. For simplicity, Figure 1 illustrates only the information for variable $x_{i}$. Suppose $m=x_{i_{1}}.x_{i_{2}}.x_{i_{3}}\ldots x_{i_{D}}$, the output of the substitution automaton $\mathcal{A}$ on the monomial $m$ is given by
@@
which is a polynomial in $\mathbb{F}[Y\sqcup Z]\mathopen{\langle}\xi\mathclose{\rangle}$. Recall that n.c. polynomials are $\mathbb{F}$-linear combinations of words/strings (called monomials). For a n.c. monomial $m$, we can identify the variable at position $c$ in $m$, where $1\leq c\leq|m|$.
@@
Let $\rho$ be a path from $q_{0}$ to $q_{s-1}$ labeled by the monomial $m$. The transformed monomial $m_{\rho}$ can be expressed in the form: $m_{\rho}=c_{m}\cdot m^{\prime}_{1}\cdot m^{\prime}_{2}\cdots m^{\prime}_{N}$, where $N\geq 1$, $c_{m}$ is a monomial over $Y\sqcup Z$, and each $m^{\prime}_{\ell}$ is given by $m^{\prime}_{\ell}=\xi^{\ell_{1}}_{1}.\xi^{\ell_{2}}_{2}\cdots\xi^{\ell_{s}}_{s}$, where $\ell_{k}>0,k\in[s-1]$ and $\ell_{s}\geq 0$.
@@
It is important to note that the number of new sub-monomials $m^{\prime}_{i}$, denoted as $N$, may not be equal to $D_{2}$. This is because $N$ depends on how many times the path $\rho$ returns to the initial state $q_{0}$. Also, the sum of exponents of $\xi$ variables in each sub-monomial $m^{\prime}_{i}$ in $m_{\rho}$ can vary. This leads us to consider two possible cases for each computation path $\rho$: (recall $m=m_{1}\cdot m_{2}\cdots m_{D_{2}}$).
@@
In Case 1, we can make the following important observation about the obtained monomial $m_{\rho}$ and the proof can be found in Appendix F. Recall that $D_{1}$ is the degree $Q_{ij}$ polynomial.
@@
Let $\rho$ be apath from $q_{0}$ to $q_{s-1}$ labeled by the monomial $m$ that satisfies Case 2. In this case, there exists a sub-monomial $m^{\prime}_{\ell}$, where $\ell\in[D_{2}]$, in the obtained monomial $m_{\rho}$ such that the sum of the exponents of its n.c. variables is not equal to $D_{1}$. That is, $\sum_{j\in[s]}\ell_{j}\neq D_{1}$.
@@
$\hat{f}=\sum\limits_{\alpha_{m,i}\cdot m\in Mon(f)}\hat{f}_{\alpha_{m,i}\cdot m}\quad\text{ and }F=\sum\limits_{\alpha_{m,i}\cdot m\in Mon(f)}F_{\alpha_{m,i}\cdot m}.$
@@
We assume that the linear forms in the $Q_{i,j}$ polynomials are numbered from 1 to $D_{1}$. For $I\subseteq[D_{1}]$ with size at most $s-1$, define $Q_{i,j,I}$ as the polynomial obtained from $Q_{i,j}$ by treating linear forms indexed by $I$ as noncommuting and the rest of the linear forms as commuting.
@@
$\hat{Q}_{ij}=\sum\limits_{I\subseteq[D_{1}],|I|=s-1}Q_{i,j,I}\times\xi_{I}.$
@@
$\hat{f}=\sum_{i\in[s]}\left(\prod_{j\in[D_{2}]}\hat{Q}_{ij}\right).$
@@
Let $f$ be a homogeneous n.c. polynomial computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$. Then, the output $f^{\prime}\in\mathbb{F}[Y\sqcup Z]\mathopen{\langle}\xi\mathclose{\rangle}$ of the substitution automaton $\mathcal{A}$ on the polynomial $f$ can expressed as $f^{\prime}=\hat{f}+F.$

### v2

@@ Abstract
In particular, we show that if $f$ is a non-zero non-commutative polynomial in $n$ variables over the field $\mathbb{F}$, computed by a depth-$d$ $+$-regular circuit of size $s$, then $f$ cannot be a polynomial identity for the matrix algebra $\mathbb{M}_{N}(\mathbb{F})$, where $N=s^{O(d^{2})}$ and the size of the field $\mathbb{F}$ depending on the degree of $f$. Interestingly, the size of the matrices does not depend on the degree of $f$. Our result can be interpreted as an Amitsur-Levitzki-type result [AL50] for polynomials computed by small-depth $+$-regular circuits.
@@ 1 Introduction
Recently, [AJMR17, AJMR19] gave an efficient randomized algorithm for the PIT problem when the circuits are allowed to compute polynomials of exponential degree, but the sparsity could be exponential in the size of the circuit. To handle doubly-exponential sparsity, they studied a class of homogeneous non-commutative circuits, that they call +-regular circuits, and gave an efficient deterministic white-box PIT algorithm. These circuits can compute non-commutative polynomials with the number of monomials doubly exponential in the circuit size. For the black-box setting, they obtain an efficient randomized PIT algorithm only for depth-3 +-regular circuits. In particular, they show that if a non-zero non-commutative polynomial $f\in\mathbb{F}\mathopen{\langle}X\mathclose{\rangle}$ is computed by a depth-3 +-regular circuit of size $s$, then $f$ cannot be a polynomial identity for the matrix algebra $\mathbb{M}_{s}(\mathbb{F})$ for a sufficiently large field $\mathbb{F}$. Finding an efficient randomized PIT algorithm for higher depth +-regular circuits is listed as an interesting open problem. We resolve this problem for constant depth +-regular circuits. In particular, we show that if $f\in\mathbb{F}\mathopen{\langle}X\mathclose{\rangle}$ is a non-zero non-commutative polynomial computed by a depth-$d$ +-regular circuit of size $s$, then $f$ cannot be a polynomial identity for the matrix algebra $\mathbb{M}_{N}(\mathbb{F})$, with $N=s^{O(d^{2})}$ and the size of the field $\mathbb{F}$ depends on the degree of polynomial $f$. This resolves an open problem given in [AJMR19]. We note that we get a black-box randomized polynomial time black-box PIT algorithm for constant depth $+$-regular circuits.
@@
The size of the $+$-regular circuit is the number of gates in the circuit. As noted earlier, the non-commutative polynomial $(x+y)^{2^{s}}$ can be computed by a depth-3 $+$-regular circuit, denoted by $\Sigma\Pi^{*}\Sigma$, of size $O(s)$ using repeated squaring. This circuit consists of two addition layers, namely $\mathcal{L}_{1}^{+},\mathcal{L}_{2}^{+}$ and two multiplication layers, namely $\mathcal{L}_{1}^{\times},\mathcal{L}_{2}^{\times}$. The multiplication layer $\mathcal{L}_{1}^{\times}$ consists of only the two input gates labeled by $x$ and $y$ respectively. The addition layer $\mathcal{L}_{1}^{+}$ consists of only one addition gate computing the homogeneous linear form $(x+y)$. The multiplication layer $\mathcal{L}_{2}^{\times}$ consists of only one gate computing the polynomial $(x+y)^{2^{s}}$. The addition layer $\mathcal{L}_{2}^{+}$ consists of only the ouptut gate computing the polynomial $(x+y)^{2^{s}}$. In this example, the top-most addition gate (i.e., the output gate) essentially has one input.
As noted in [AJMR19], the computational power of +-regular circuits is limited. Nisan’s rank-based argument [Nis91] can be adapted to show that the non-commutative permanent cannot be computed by polynomial-size +-regular circuits. However, despite their structural constraints, polynomial-size +-regular circuits with +-depth 2 can compute polynomials with exponential degree and a doubly exponential number of monomials.
The main result of the paper is the following theorem.
@@
For degree $D$ non-zero non-commutative polynomial $f$, the classical Amitsur-Levitzki [AL50] theorem guarantees that $f$ does not vanish on the matrix algebra $\mathcal{M}_{\frac{D}{2}+1}(\mathbb{F})$. If $D=2^{\Omega(s)}$, this gives us an exponential time randomized PIT algorithm [BW05], where $s$ is the size of the circuit computing $f$. In contrast, in our result, the dimension of the matrices is independent of the degree of the polynomial. If the sparsity of the polynomial, i.e., the number of non-zero monomials, is doubly exponential, then the main result of [AJMR19] gives only an exponential time randomized PIT algorithm as their matrix dimension depends on the logarithm of the sparsity.
@@
We note that [AJMR19] presented a white-box deterministic polynomial-time PIT for arbitrary depth $+$-regular circuits. For the small-degree case, [RS05] provided a white-box deterministic polynomial-time PIT for non-commutative ABPs, while [FS13, AGKS15] have shown a deterministic quasi-polynomial-time black-box PIT algorithm for non-commutative ABPs. In the commutative setting, a randomized polynomial time PIT was given by [IM83].
@@ 1.1 Outline of the Proofs: High-level Idea
In the rest of the paper, we will use ”n.c.” as an abbreviation for ”non-commutative.” We begin by explaining the main ideas behind the randomized PIT algorithm for depth-5 $+$-regular circuits, as this represents the primary bottleneck. The existing techniques are not directly applicable to this case, and solving the depth-5 case requires several new ideas, which we outline below. It turns out that this is the key bottleneck, as we can later adapt the results from the depth-5 $+$-regular circuits to higher depths with additional techniques.
Consider an n.c. polynomial $f$ over $X=\{x_{1},\ldots,x_{n}\}$ computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$. The polynomial $f$, computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$, can be written as follows:
@@ 1.1.1 Step 1: Transforming the Polynomial for Improved Structure
This fact is formalized in Lemma 4.1.1. However, this fact alone will not be sufficient to design an identity-testing algorithm for depth-5 +-regular circuits. This is because, to choose a small index set from each $Q_{ij}$, one would have to know the boundary between $Q_{ij}$ and $Q_{i,j+1}$ (i.e., the position at which $Q_{ij}$ ends and $Q_{i,j+1}$ begins), for any $i\in[s],j\in[D_{2}-1]$.
Definition 1.1.1 (Boundary of a Polynomial).
Let $P=\prod_{i\in[k]}P_{i}$ be a non-commutative polynomial over $X$. Each non-zero monomial $m$ of $P$ can be expressed as: $m=m_{1}m_{2}\cdots m_{k}$. We say the boundary of $P_{i}$, $i\in[k]$ as the position at which $P_{i}$ ends. Similarly, the boundary of $m_{i}$, $i\in[k]$ in the monomial $m$ is defined as the position at which $m_{i}$ ends.
Since non-commutative polynomials are $\mathbb{F}$-linear combinations of words/strings (called monomials) over $X$, for a n.c. monomial $m$, we can identify the variable at position $e$ in $m$, where $1\leq e\leq|m|$.
To know where $Q_{ij}$ ends and $Q_{i,j+1}$ begins exactly, the algorithm will have to keep track of the degree (length) of $Q_{ij}$. This is not feasible as the degree of $Q_{ij}$ could be exponential in $s$. This is one of the main challenges in designing a black-box identity-testing algorithm for depth-5 +-regular circuits and is one of the main differences between depth-3 and depth-5 +-regular circuits (small depth $+$-regular circuits, in general).
@@ 1.1.5 Matrix Compositions
Although [AJMR19] provides a randomized polynomial-time black-box PIT algorithm for $+$-regular circuits of depth 3, existing techniques are not directly applicable to this case. In particular, depth-5 $+$-regular circuits require new ideas, which we introduce in our algorithm. While our algorithm operates in a black-box setting, it can be described as a sequence of transformations that convert the non-commutative polynomial into a commutative one, enabling identity testing. In Lemma 3.0.1, we show that the matrix substitutions can be composed into a single matrix for each input variable, capturing the entire transformation process. Our coefficient modification approach introduces a novel technique in polynomial identity testing, which may be useful for developing PIT algorithms for other circuit models.
@@ 2.1 Substitution Automaton
We observe that all the automatons we design are more restrictive in that, for any $i\in Q$ and $x\in X$, if both $(j,u)$ and $(j,v)$ belong to $\psi(i,x)$, then $u$ must equal $v$. In other words, the substitution applied to variable $x$ when transitioning from state $i$ to state $j$ is uniquely determined by the destination state $j$.
@@
Now, we explain how a substitution automaton $\mathcal{A}$ processes a given polynomial $f$. First, we will describe the output of a substitution automaton $\mathcal{A}$ on a monomial $w\in X^{D}$. Let $s$ and $t$ be the designated initial and final states of $\mathcal{A}$, respectively. For each variable $x\in X$, we have a transition matrix $M^{\prime}_{x}$ of size $|Q|\times|Q|$. If $w=x_{j_{1}}x_{j_{2}}\ldots x_{j_{d}}$, then the output of the substitution automaton $\mathcal{A}$ on the monomial $w$ is given by the $(s,t)$-th entry of the matrix $M^{\prime}_{w}=M^{\prime}_{x_{j_{1}}}M^{\prime}_{x_{j_{2}}}\ldots M^{\prime}_{x_{j_{D}}}$. Note that this entry can be a polynomial since there could be multiple paths from $s$ to $t$ labeled by the same monomial $w$. The output of the substitution automaton $\mathcal{A}$ on the given polynomial $f$ is provided by the $(s,t)$-th entry of the matrix $f(M^{\prime}_{x_{1}},M^{\prime}_{x_{2}},\ldots,M^{\prime}_{x_{n}})$, which is obtained by substituting the matrix $M^{\prime}_{x_{i}}$ for each variable $x_{i}$ where $i\in[n]$ in the polynomial.
The substitution automaton processes each monomial of $f$ and sums over the resulting monomials.
@@ 4 Black-Box PIT for $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ Circuits
Definition 4.0.2 ($\xi$-Pattern/Ordered Power-Sum Monomial).
We refer to the string $\xi_{1}^{\ell_{1}}\xi_{2}^{\ell_{2}}\cdots\xi_{k}^{\ell_{k}}$ as a $\xi$-pattern, regardless of the specific exponents of the $\xi$ variables. We also refer $\xi$-pattern as an ordered power-sum monomial.
@@ 4.1.1 Idea of the Substitution Automaton
    As explained in the proof outline, one would have to identify the boundary between $Q_{ij}$ and $Q_{i,j+1}$ (i.e., the position at which $Q_{ij}$ ends and $Q_{i,j+1}$ begins), for any $i\in[s],j\in[D_{2}-1]$. This task was straightforward in [AJMR19] because each $Q_{ij}$ is a linear form in their work.
@@
    We need to ensure that the resulting commutative polynomial is non-zero because the same commutative variables appearing in different transformed $Q_{ij}$ polynomials get mixed (i.e., exponents of the same variable appearing at different transformed $Q_{ij}$ polynomials are added). Additionally, after converting to the commutative polynomial, we must guarantee that the commutative counterpart of each $Q_{ij}$ polynomial is non-zero. In the case of [AJMR19], since each $Q_{ij}$ is a linear form, we can treat them as commutative without the concern that $Q_{ij}$ may become a zero polynomial. However, as noted, if the degree of $Q_{ij}$ is greater than 1, we cannot guarantee non-zero values by merely considering them commutative. For example, the non-commutative polynomial $xy-yx$ becomes zero if $x$ and $y$ are allowed to commute.
@@
which is a polynomial in the variables $\xi\sqcup Y\sqcup Z$.
@@
Next, we consider the output of the substitution automaton $\mathcal{A}$ on $m$. The automaton knows how to replace/substitute any variable $x_{j}$ at any state $q$. For simplicity, Figure 1 illustrates only the information for variable $x_{i}$. Suppose $m=x_{i_{1}}.x_{i_{2}}.x_{i_{3}}\ldots x_{i_{D}}$, the output of the substitution automaton $\mathcal{A}$ on the monomial $m$ is given by
@@
which is a polynomial in $\mathbb{F}[Y\sqcup Z]\mathopen{\langle}\xi\mathclose{\rangle}$. Recall that n.c. polynomials are $\mathbb{F}$-linear combinations of words/strings (called monomials). For a n.c. monomial $m$, we can identify the variable at position $e$ in $m$, where $1\leq e\leq|m|$.
@@
Let $\rho$ be a path from $q_{0}$ to $q_{s-1}$ labeled by the monomial $m$. The transformed monomial $m_{\rho}$ can be expressed in the form: $m_{\rho}=c_{m}\cdot m^{\prime}_{1}\cdot m^{\prime}_{2}\cdots m^{\prime}_{N}$, where $N\geq 1$, $c_{m}$ is a monomial over $Y\sqcup Z$, and each $m^{\prime}_{\ell}$ (for $\ell\in[N]$) is given by $m^{\prime}_{\ell}=\xi^{\ell_{1}}_{1}.\xi^{\ell_{2}}_{2}\cdots\xi^{\ell_{s}}_{s}$, where $\ell_{k}>0$ for $k\in[s-1]$, and $\ell_{s}\geq 0$.
@@
It is important to note that the number of new sub-monomials $m^{\prime}_{i}$, denoted as $N$, may not be equal to $D_{2}$. This is because $N$ depends on how many times the path $\rho$ returns to the initial state $q_{0}$ (see Figure 1). Also, the sum of exponents of $\xi$ variables in each sub-monomial $m^{\prime}_{i}$ in $m_{\rho}$ can vary. This leads us to consider two possible cases for each computation path $\rho$ that starts at $q_{0}$ and ends at $q_{s-1}$: (recall $m=m_{1}\cdot m_{2}\cdots m_{D_{2}}$).
@@
In Case 1, we can make the following important observation about the obtained monomial $m_{\rho}$ and the proof can be found in Appendix F. Recall that $D_{1}$ is the degree of $Q_{ij}$ polynomial for all $i\in[s]$ and $j\in[D_{2}]$.
@@
Let $\rho$ be apath from $q_{0}$ to $q_{s-1}$ labeled by the monomial $m$ that satisfies Case 2. In this case, there exists a sub-monomial $m^{\prime}_{\ell}$, where $\ell\in[N]$, in the obtained monomial $m_{\rho}$ such that the sum of the exponents of its n.c. variables is not equal to $D_{1}$. That is, $\sum_{j\in[s]}\ell_{j}\neq D_{1}$.
@@
$\hat{f}=\sum\limits_{\alpha_{m,i}\cdot m\in Mon(f)}\hat{f}_{\alpha_{m,i}\cdot m}$ (4)
$F=\sum\limits_{\alpha_{m,i}\cdot m\in Mon(f)}F_{\alpha_{m,i}\cdot m}.$ (5)
@@
We assume that the linear forms in the $Q_{i,j}$ polynomials are numbered from 1 to $D_{1}$. For $I\subseteq[D_{1}]$ with size at most $s-1$, define $Q_{i,j,I}$ as the polynomial obtained from $Q_{i,j}$ by treating linear forms indexed by $I$ as non-commuting and the rest of the linear forms as commuting. We also substitute the n.c variables that appear in linear forms indexed by $I$ with double-indexed commutative variables $Y$, as shown in the substitution automaton (see Figure 1).
@@
$\hat{Q}_{ij}=\sum\limits_{I\subseteq[D_{1}],|I|=s-1}Q_{i,j,I}\times\xi_{I}.$ (6)
@@
$\hat{f}=\sum_{i\in[s]}\left(\prod_{j\in[D_{2}]}\hat{Q}_{ij}\right).$ (7)
@@
Let $f$ be a homogeneous n.c. polynomial computed by a $\Sigma\Pi^{*}\Sigma\Pi^{*}\Sigma$ circuit of size $s$. Then, the output $f^{\prime}\in\mathbb{F}[Y\sqcup Z]\mathopen{\langle}\xi\mathclose{\rangle}$ of the substitution automaton $\mathcal{A}$ on the polynomial $f$ can expressed as $f^{\prime}=\hat{f}+F,$ where $\hat{f}$ be the structured and spurious part as defined in Equation 7 and $F$ be the spurious part as defined in Equation 5.